# On Jiang’s wavefront sets conjecture for representations in local Arthur packets

### 13 Mar 2026 v1

@@ 4.2. Dimension identities for nilpotent orbits
We have the following dimension identity for nilpotent orbits.
@@
(4.8) $\begin{split}&\,\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}([b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]))\\ =&\,\dim(\hat{\mathfrak{g}}_{1})+\dim(\hat{\mathfrak{g}}_{2})-\dim_{\hat{\mathfrak{g}}_{1}}(([\prod_{i\in I}b_{i}^{a_{i}}]^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}})-\dim_{\hat{\mathfrak{g}}_{2}}(([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).\end{split}$
@@ 5. Construction of elements in local Arthur packets
In this section, we prove Theorems 1.9 and 1.13. We follow closely the notation in Section 4.
First we recall the identities (4.3), (4.5), and (4.7). When ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, the identity (4.3) is as follows,
@@
(6.1) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
@@
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, transferring to $f^{\prime}=f^{\prime 1}\otimes f^{\prime 2}\in\widetilde{{\mathcal{S}}}(G^{\prime})$, $f^{\prime i}\in\widetilde{{\mathcal{S}}}(G^{\prime}_{i})$, $i=1,2$. Let $\widetilde{f}^{1}\in\widetilde{{\mathcal{H}}}(2n_{1}+1)$, and $\widetilde{f}^{2}\in\widetilde{{\mathcal{H}}}(2n_{2})$, transferring to $f^{\prime 1}$ and $f^{\prime 2}$, via surjective maps in [Art13, Corollary 2.1.2]
$\iota_{G_{k}}:\widetilde{{\mathcal{H}}}(N)\rightarrow\widetilde{{\mathcal{S}}}(G_{k}),k=n_{1},n_{2},$
respectively. Here, $\psi^{1}=\sum_{i\in I}\psi_{i}$, $\psi^{2}=\sum_{j\in J}\psi_{j}$, with $I=\{i|b_{i}\text{ odd}\}$, $J=\{i|b_{i}\text{ even}\}$. $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ is defined in (1.3), and $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}^{i}$ is the Harish-Chandra descent of $\widetilde{f}^{i}$ (see for example [Kon02, Section 3.1]), $i=1,2$. $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)$, where $c_{{\mathrm{G}}_{n}}=-I_{2n}$. By Conjecture 1.2 and Remark 1.4 Part (3), the maximal nilpotent orbits occurring on the right hand side of (6.1) are those corresponding to the partitions
$(([\prod_{i\in I}b_{i}^{a_{i}}]^{t})^{-})_{{\mathrm{G}}_{1}^{\prime}}=(([\prod_{i\in I}b_{i}^{a_{i}}]^{-})_{{\mathrm{G}}_{1}^{\prime}})^{t}=\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1}))$
and $([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$.
When ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, the identities (4.5) and (4.7) are as follows,
(6.2) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, transferring to $f^{\prime}=f^{\prime 1}\otimes f^{\prime 2}\in\widetilde{{\mathcal{S}}}(G^{\prime})$, $f^{\prime i}\in\widetilde{{\mathcal{S}}}(G^{\prime}_{i})$, $i=1,2$. Let $\widetilde{f}^{1}\in\widetilde{{\mathcal{H}}}(2n_{1})$ and $\widetilde{f}^{2}\in\widetilde{{\mathcal{H}}}(2n_{2})$, transferring to $f^{\prime 1}$ and $f^{\prime 2}$, via surjective maps in [Art13, Corollary 2.1.2]
$\iota_{G_{k}}:\widetilde{{\mathcal{H}}}(N)\rightarrow\widetilde{{\mathcal{S}}}(G_{k}),k=n_{1},n_{2},$
@@
respectively. Here, $\psi^{1}=\sum_{i\in I}\psi_{i}$, $\psi^{2}=\sum_{j\in J}\psi_{j}$, with $I=\{i|b_{i}\text{ odd}\}$, $J=\{i|b_{i}\text{ even}\}$, $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ is defined in (1.3), and $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}^{i}$ is the Harish-Chandra descent of $\widetilde{f}^{i}$ (see for example [Kon02, Section 3.1]), $i=1,2$. $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)$, where
$c_{{\mathrm{G}}_{n}}=\begin{cases}I_{2n+1}&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},\\ -I_{2n}&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}.\end{cases}$
By Conjecture 1.2, the maximal nilpotent orbits occurring on the right hand side of (6.2) are those corresponding to the partitions $([\prod_{i\in I}b_{i}^{a_{i}}]^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}$ and $([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$.
@@
We next recall several lemmas from [Sha90] and [Kon02] as follows. Let $\Delta$ be the transfer factor defined by Langlands-Shelstad ([LS87]), or by Kottwitz-Shelstad ([KS99]).
@@
[Kon02, Lemma 8.5] Suppose $\widetilde{f}\in\widetilde{\mathcal{H}}(N)$ has sufficiently small support near $1$. Then for $t\in F^{\times}$ with $|t|$ small, $\widetilde{f}_{t^{2}}$ and $(\widetilde{f}^{G_{n}})_{t^{2}}$ have $\Delta$-twisted matching orbital integrals.
Now, we are ready to prove Theorem 1.9. First, we prove Part (1) of Theorem 1.9.
@@
Let $\psi$ be a local Arthur parameter as in (1.1). Assume that Conjecture 1.2 holds, then Conjecture 1.7 Part (2) is valid.
@@
Assume that there there exist $\underline{p}>\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))$ and $\pi\in\widetilde{\Pi}_{\psi}$ such that $\underline{p}\in\mathfrak{p}^{m}(\pi)$. Without loss of generality, we assume further that

### 06 May 2026 v2

@@ 4.2. Dimension identities for nilpotent orbits
We first record a dimension identity as follows.
@@
(4.4) $\begin{split}&\,\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}([b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]))\\ =&\,\dim(\hat{\mathfrak{g}}^{\prime}_{1})+\dim(\hat{\mathfrak{g}}^{\prime}_{2})-\dim_{\hat{\mathfrak{g}}^{\prime}_{1}}(([\prod_{i\in I}b_{i}^{a_{i}}]^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}})-\dim_{\hat{\mathfrak{g}}^{\prime}_{2}}(([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).\end{split}$
@@
begin
Remark 4.3.
In our setting, Lemma 4.1 must be replaced by corresponding analogues in which the right hand side is computed in the Lie algebras of the connected stabilizers ${{\mathrm{G}}_{i}^{\prime}}^{\theta}$ of the points $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}=1\rtimes\widetilde{\theta}(N_{i})$. Concretely, the desired target partitions are
${(\underline{p}(\psi^{1})^{t})_{{\mathrm{SO}}_{2n_{1}+1}}\text{ and }(\underline{p}(\psi^{2})^{t})_{{\mathrm{Sp}}_{2n_{2}}},}$ ${\qquad\text{when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},}$
${(\underline{p}(\psi^{i})^{t})_{{\mathrm{Sp}}_{2n_{i}}},\ i=1,2,}$ ${\qquad\text{when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha}.}$
Proposition 4.4.
The following dimension identity holds:
$\begin{split}&{\,\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}(}[{b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}}]{))}\\ {=}&{\,\dim(\mathfrak{g_{1}^{\prime}}^{\theta})+\dim(\mathfrak{g_{2}^{\prime}}^{\theta})-\dim_{\mathfrak{g_{1}^{\prime}}^{\theta}}((}[{\prod_{i\in I}b_{i}^{a_{i}}}]{{}^{t})_{{{\mathrm{G}}^{\prime}_{1}}^{\theta}})-\dim_{\mathfrak{g_{2}^{\prime}}^{\theta}}((}[{\prod_{j\in J}b_{j}^{a_{j}}}]{{}^{t})_{{{\mathrm{G}}^{\prime}_{2}}^{\theta}}).}\end{split}{}$
Proof.
The split special odd orthogonal case is already compatible with Lemma 4.1. The symplectic and special even orthogonal cases can be obtained by a simple codimension comparison as follows.
Assume that $\mathrm{G}_{n}={\mathrm{Sp}}_{2n}$. By Lemma 4.1, it is enough to replace the even orthogonal contribution on the right hand side by the corresponding symplectic contribution. Let
${\underline{p}_{J}=}[{\prod_{j\in J}b_{j}^{a_{j}}}]{{}^{t}.}$
Since $\mathrm{G}_{n}={\mathrm{Sp}}_{2n}$, the partition $b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]$ is orthogonal. Hence every even part $b_{j}$ occurs with even multiplicity $a_{j}$, $j\in J$. It follows that every part of $\underline{p}_{J}$ is even, because the parts of $\underline{p}_{J}$ are partial sums of the even integers $a_{j}$. Moreover each part of $\underline{p}_{J}$ occurs with even multiplicity, since these multiplicities are differences of even parts $b_{j}$ (with 0 allowed at the end). Therefore, $\underline{p}_{J}$ is simultaneously an orthogonal partition and a symplectic partition of $2n_{2}$, and
$(\underline{p}_{J})_{{\mathrm{SO}}_{2n_{2}}}=(\underline{p}_{J})_{{\mathrm{Sp}}_{2n_{2}}}=\underline{p}_{J}.$}
Writing $\underline{p}_{J}=[q_{1}q_{2}\cdots q_{l}]$ with $q_{1}\geq\cdots\geq q_{l}$ and using the notation $r_{i}^{\underline{p}_{J}}$ and $s_{i}^{\underline{p}_{J}}$ from the proof of Lemma 4.1, we have
$\sum_{i\text{ odd }}r_{i}^{\underline{p}_{J}}=0,$}
because all parts of $\underline{p}_{J}$ are even. Hence, by [CM93, Corollary 6.1.4],
${\dim(\mathfrak{sp}_{2n_{2}})-\dim_{\mathfrak{sp}_{2n_{2}}}(\underline{p}_{J})}$ ${=\frac{1}{2}\sum_{i}\left(s_{i}^{\underline{p}_{J}}\right)^{2}}$
${=\dim(\mathfrak{so}_{2n_{2}})-\dim_{\mathfrak{so}_{2n_{2}}}(\underline{p}_{J}).}$
Thus the even-factor codimension appearing in Lemma 4.1 is unchanged when we pass from $\mathfrak{so}_{2n_{2}}$ to $\mathfrak{sp}_{2n_{2}}$, and the displayed identity follows.
Next we assume that $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$. By Lemma 4.1, it is enough to replace the two orthogonal factor codimensions on its right hand side by the corresponding symplectic factor codimensions. Put
$\underline{p}_{I}=[\prod_{i\in I}b_{i}^{a_{i}}]^{t},\qquad\underline{p}_{J}=[\prod_{j\in J}b_{j}^{a_{j}}]^{t}.$}
If one of $I,J$ is empty, the corresponding rank-zero assertion is understood in the evident sense.
The $J$-factor is exactly the same comparison as in the ${\mathrm{Sp}}_{2n}$ case above. Since $b_{1}^{a_{1}}\cdots b_{r}^{a_{r}}]$ is an orthogonal partition, every even $b_{j}$ has even multiplicity. Hence every part of $\underline{p}_{J}$ is even, and every part occurs with even multiplicity. Thus
[…]
Using the dimension formulas in [CM93, Corollary 6.1.4], this gives
$\dim(\mathfrak{so}_{2n_{2}})-\dim_{\mathfrak{so}_{2n_{2}}}(\underline{p}_{J})=\frac{1}{2}\sum_{a}(s_{a}^{\underline{p}_{J}})^{2}=\dim(\mathfrak{sp}_{2n_{2}})-\dim_{\mathfrak{sp}_{2n_{2}}}(\underline{p}_{J}).$}
It remains to compare the $I$-factor. Write
[…]
and set $C_{k}=c_{1}+\cdots+c_{k}$. Since the total size $\sum_{k}c_{k}d_{k}$ is even and all $d_{k}$ are odd, $C_{t}$ is even. The transpose has the form
$\underline{p}_{I}=[C_{t}^{d_{t}}C_{t-1}^{d_{t-1}-d_{t}}\cdots C_{1}^{d_{1}-d_{2}}].$}
For $k<t$, the multiplicity $d_{k}-d_{k+1}$ is even. Therefore every odd part of $\underline{p}_{I}$ has even multiplicity, and $\underline{p}_{I}$ is a symplectic partition; in particular $(\underline{p}_{I})_{{\mathrm{Sp}}_{2n_{1}}}=\underline{p}_{I}$. Let
$\mu_{I}=(\underline{p}_{I})_{{\mathrm{SO}}_{2n_{1}}}.$}
The $D$-collapse recipe of [CM93, Lemma 6.3.8], applied to the displayed form of $\underline{p}_{I}$, gives
$\sum_{a}(s_{a}^{\mu_{I}})^{2}-\sum_{a}(s_{a}^{\underline{p}_{I}})^{2}=\sum_{a\text{ odd}}r_{a}^{\mu_{I}}+\sum_{a\text{ odd}}r_{a}^{\underline{p}_{I}}.$}
Substituting this equality into the symplectic and even orthogonal dimension formulas in [CM93, Corollary 6.1.4] yields
$\dim(\mathfrak{sp}_{2n_{1}})-\dim_{\mathfrak{sp}_{2n_{1}}}(\underline{p}_{I})=\dim(\mathfrak{so}_{2n_{1}})-\dim_{\mathfrak{so}_{2n_{1}}}(\mu_{I}).$}
Combining the $I$- and $J$-factor codimension equalities with Lemma 4.1 gives the displayed identity. ∎
At the end of this section, we record another dimension equality which will be used in later sections.
Lemma 4.5.
Let $\lambda$ be an orthogonal partition of $2n$, and put $\tau=\lambda^{t}$. Then
$\dim(\mathfrak{so}_{2n})-\dim_{\mathfrak{so}_{2n}}(\tau_{{\mathrm{SO}}_{2n}})=\dim(\mathfrak{sp}_{2n})-\dim_{\mathfrak{sp}_{2n}}(\tau_{{\mathrm{Sp}}_{2n}}).$}
Proof.
Set
$\rho=\tau_{{\mathrm{Sp}}_{2n}},\qquad\mu=\tau_{{\mathrm{SO}}_{2n}}.$}
We first recall the compatibility of the two collapses in this special situation. Since $\lambda=\tau^{t}$ is an orthogonal partition, the explicit collapse recipe of [CM93, Lemma 6.3.8] gives
$\mu=(\tau_{{\mathrm{Sp}}_{2n}})_{{\mathrm{SO}}_{2n}}=\rho_{{\mathrm{SO}}_{2n}},$}
and $\rho^{t}$ is again an orthogonal partition. Indeed, writing $r_{a}^{\tau}=\lambda_{a}-\lambda_{a+1}$, the parts of $\tau$ which are bad for the $C$-collapse are exactly the odd integers $a$ for which the two consecutive column lengths $\lambda_{a}$ and $\lambda_{a+1}$ have opposite parity. Let $u,\ldots,v$ be a maximal block of even entries in the ordered partition $\lambda$. Its length is even. Hence the two boundary indices u1u-1 and $v$ (when both boundaries occur) have the same parity. If this common parity is odd, the $C$-collapse pairs the two bad odd parts $v$ and u1u-1 and replaces them by the adjacent even parts v1v-1 and $u$; if this common parity is even, no $C$-move is made at that block and the two boundary parts remain bad for the subsequent $D$-collapse. At the top or bottom of $\lambda$ there is only one boundary, and its index is even, so it is also a $D$-boundary. Thus the $C$-collapse only removes the odd bad boundary pairs coming from even blocks of $\lambda$, and after these replacements the even column-length blocks still have even length. This is the assertion that $\rho^{t}$ is orthogonal. The remaining bad even parts, with the usual terminal 0 if needed, are then exactly the $D$-collapse chains which give the direct $D$-collapse of $\tau$; hence the displayed identity for $\mu$.
It remains to prove the codimension identity for such a partition $\rho$. We write $r_{a}^{\sigma}$ and $s_{a}^{\sigma}$ for the multiplicities and column lengths of any partition $\sigma$, as in the proof of Lemma 4.1. By [CM93, Corollary 6.1.4], for a symplectic partition $\sigma$ of $2n$ and an orthogonal partition $\nu$ of $2n$,
${\mathrm{cod}}_{\mathfrak{sp}_{2n}}(\sigma)=\frac{1}{2}\left(\sum_{a}(s_{a}^{\sigma})^{2}+\sum_{a\text{ odd}}r_{a}^{\sigma}\right),$}
${\mathrm{cod}}_{\mathfrak{so}_{2n}}(\nu)=\frac{1}{2}\left(\sum_{a}(s_{a}^{\nu})^{2}-\sum_{a\text{ odd}}r_{a}^{\nu}\right),$}
where ${\mathrm{cod}}_{\mathfrak{h}}$ means $\dim(\mathfrak{h})$ minus the orbit dimension. We shall prove
${\mathrm{cod}}_{\mathfrak{so}_{2n}}(\rho_{{\mathrm{SO}}_{2n}})={\mathrm{cod}}_{\mathfrak{sp}_{2n}}(\rho).$}
Let $\nu=\rho_{{\mathrm{SO}}_{2n}}$. Since $\rho$ is symplectic, its odd parts occur with even multiplicity; since $\rho^{t}$ is orthogonal, every odd part of $\rho$ lies between two ends of one of the $D$-collapse chains below. Group the even parts of $\rho$ which have odd multiplicity into the standard $D$-collapse chains. One such chain has even entries
$2d_{0}>2d_{1}>\cdots>2d_{h}\geq 0,$}
where $2d_{0}$ and $2d_{h}$ have odd multiplicity (with $2d_{h}=0$ allowed as the terminal dummy part), and $2d_{1},\ldots,2d_{h-1}$ are the even parts of $\rho$ strictly between them. On this chain the $D$-collapse replaces one copy of $2d_{0}$ by $2d_{0}-1$, one copy of $2d_{h}$ by $2d_{h}+1$ if $d_{h}>0$ (and adds one copy of $1$ if $d_{h}=0$), and for each intermediate even part $2d_{j}$ replaces two copies of $2d_{j}$ by $2d_{j}+1$ and $2d_{j}-1$. Odd parts already present in $\rho$ are left unchanged.
For this chain, the changes in the column lengths are exactly as follows: $s_{2d_{0}}$ decreases by $1$, $s_{2d_{h}+1}$ increases by $1$, and for each intermediate $j$ both $s_{2d_{j}}$ decreases by $1$ and $s_{2d_{j}+1}$ increases by $1$. Denote by $\Delta_{D}$ the contribution of this single $D$-collapse chain to the total change $\sum_{a}(s_{a}^{\nu})^{2}-\sum_{a}(s_{a}^{\rho})^{2}$. Then
$\Delta_{D}=-2s_{2d_{0}}^{\rho}+1+2s_{2d_{h}+1}^{\rho}+1\;+\sum_{j=1}^{h-1}\left(2(s_{2d_{j}+1}^{\rho}-s_{2d_{j}}^{\rho})+2\right).$}
Put
$I_{D}=\{a\mid a\text{ is odd and }2d_{h}<a<2d_{0}\}.$}
Since $s_{2d_{h}+1}^{\rho}-s_{2d_{0}}^{\rho}$ counts all parts of $\rho$ with sizes in the interval $2d_{h}+1,2d_{0}-1]$, and since $s_{2d_{j}+1}^{\rho}-s_{2d_{j}}^{\rho}=-r_{2d_{j}}^{\rho}$, we get the exact cancellation
$s_{2d_{h}+1}^{\rho}-s_{2d_{0}}^{\rho}+\sum_{j=1}^{h-1}(s_{2d_{j}+1}^{\rho}-s_{2d_{j}}^{\rho})=\sum_{a\in I_{D}}r_{a}^{\rho}.$}
Thus this single-chain contribution is
$\Delta_{D}=2\sum_{a\in I_{D}}r_{a}^{\rho}+2h.$}
Now compare this with the odd multiplicities. The $D$-collapse does not change any old odd part, and the only odd parts created by this chain lie in $I_{D}$: one from $2d_{0}$, one from $2d_{h}$ (with $2d_{h}=0$ contributing the new part $1$), and two from each intermediate $2d_{j}$. Counted with multiplicity, this creates exactly $2h$ odd parts in $I_{D}$. Hence
$\sum_{a\in I_{D}}r_{a}^{\nu}=\sum_{a\in I_{D}}r_{a}^{\rho}+2h,$}
and therefore
$\Delta_{D}=\sum_{a\in I_{D}}r_{a}^{\nu}+\sum_{a\in I_{D}}r_{a}^{\rho}.$}
The intervals $I_{D}$ for the distinct collapse chains are disjoint and together account for exactly the odd parts whose multiplicities enter the change. Summing over the chains gives
$\sum_{a}(s_{a}^{\nu})^{2}-\sum_{a}(s_{a}^{\rho})^{2}=\sum_{a\text{ odd}}r_{a}^{\nu}+\sum_{a\text{ odd}}r_{a}^{\rho}.$}
Substituting this identity into the two codimension formulas above proves
${\mathrm{cod}}_{\mathfrak{so}_{2n}}(\rho_{{\mathrm{SO}}_{2n}})={\mathrm{cod}}_{\mathfrak{sp}_{2n}}(\rho).$}
Since $\rho=\tau_{{\mathrm{Sp}}_{2n}}$ and $\rho_{{\mathrm{SO}}_{2n}}=\tau_{{\mathrm{SO}}_{2n}}$, this is the desired identity. ∎
end
@@ 5. Construction of elements in local Arthur packets
In this section, we prove Theorems 1.9. We follow closely the notation in Section 4. First we recall the identity (4.3).
@@
(6.1) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(f)\\ =&\,\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
@@
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, transferring to $f^{\prime}=f^{\prime 1}\otimes f^{\prime 2}\in\widetilde{{\mathcal{S}}}(G^{\prime})$, $f^{\prime i}\in\widetilde{{\mathcal{S}}}(G^{\prime}_{i})$, $i=1,2$. Let $\widetilde{f}^{1}\in\widetilde{{\mathcal{H}}}(N_{1})$, and $\widetilde{f}^{2}\in\widetilde{{\mathcal{H}}}(N_{2})$, transferring to $f^{\prime 1}$ and $f^{\prime 2}$, via surjective maps in [Art13, Corollary 2.1.2]
$\iota_{G^{\prime}_{i}}:\widetilde{{\mathcal{H}}}(N_{i})\rightarrow\widetilde{{\mathcal{S}}}(G^{\prime}_{i}),i=1,2,$
@@
respectively. Here, $\psi^{1}=\sum_{i\in I}\psi_{i}$, $\psi^{2}=\sum_{j\in J}\psi_{j}$, with $I=\{i|b_{i}\text{ odd}\}$, $J=\{i|b_{i}\text{ even}\}$. $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ is defined in (1.3), and $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}^{i}$ is the Harish-Chandra descent of $\widetilde{f}^{i}$ (see for example [Kon02, Section 3.1]), $i=1,2$. $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}c_{{\mathcal{O}}}(\pi)$.
@@
Next, we recall several lemmas from [Sha90] and [Kon02] as follows. Let $\Delta$ be the transfer factor defined by Langlands-Shelstad ([LS87]), or by Kottwitz-Shelstad ([KS99]).
@@
[Kon02, Lemma 8.5] Suppose $\widetilde{f}\in\widetilde{\mathcal{H}}(N)$ has sufficiently small support near $1$ and transfers to $f^{G_{n}}$ for some $f\in\widetilde{\mathcal{H}}(G_{n})$. Then for $t\in F^{\times}$ with $|t|$ small, ${f}^{G_{n}}_{t^{2}}$ and $|t|^{\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g}_{n}^{\theta})}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}})_{t^{2}}$ have $\Delta$-twisted matching orbital integrals.
The above statement in Lemma 6.3 is directly implied by the same proof of [Kon02, Lemma 8.5], while [Kon02, Lemma 8.5] considered the setting of $\dim(\mathfrak{g}_{n})=\dim(\mathfrak{g}_{n}^{\theta})$.
Now, we are ready to prove Theorem 1.9. First, we prove 1.9(1).
@@
Let $\psi$ be a local Arthur parameter as in (1.1). Then, Conjecture 1.2 is true if and only if for any $\pi\in\widetilde{\Pi}_{\psi}$, $\underline{p}\in\mathfrak{p}^{m}(\pi)$,
$\dim_{{\mathfrak{g}}_{n}}(\underline{p})\leq\dim_{{\mathfrak{g}}_{n}}(\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))).$
@@
First, we assume that Conjecture 1.2 is true. Assume that there there exist $\pi\in\widetilde{\Pi}_{\psi}$ and $\underline{p}\in\mathfrak{p}^{m}(\pi)$, such that $\dim_{{\mathfrak{g}}_{n}}(\underline{p})>\dim_{{\mathfrak{g}}_{n}}(\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi)))$. Without loss of generality, we assume further that